# A Dataset and Benchmark for Hospital Course Summarization with Adapted Large Language Models

### v2

@@ Abstract
Materials and Methods: Using clinical notes as input, we apply prompting-based (using in-context learning) and fine-tuning-based adaptation strategies to three open-source LLMs (Clinical-T5-Large, Llama2-13B, FLAN-UL2) and two proprietary LLMs (GPT-3.5, GPT-4). We evaluate these LLMs across multiple context-length inputs using natural language similarity metrics. We further conduct a clinical study with five clinicians, comparing clinician-written and LLM-generated BHCs across 30 samples, focusing on their potential to enhance clinical decision-making through improved summary quality. We compare reader preferences for the original and LLM-generated summary using Wilcoxon Signed-Rank tests.
@@
Clinicians spend significant time on clinical documentation [moy2021measurement, chaiyachati2019assessment, mamykina2016residents], a crucial component of data assimilation [albers2020assimilation] beginning with data sources such as EHRs and progressing through data translation and knowledge creation where these records can be processed for clinically-relevant machine learning tasks. The discharge summary (clinical note) is a key document clinicians refer to in important scenarios, including when a discharged patient presents to the clinic for post-hospitalization follow-up. The brief hospital course (BHC) fulfills an important role within the discharge summary, particularly for long and complex hospitalizations. Clinicians who are pressed for time may choose to focus on the BHC summary instead of the entire discharge summary. Synthesizing a BHC given a clinical note requires substantial clinician time and expertise, with errors possibly causing patient harm [clough2024transforming]. Automating BHC generation is important since discharge summaries can lack essential information, contain incorrect information, and may be incomplete before follow-up appointments [kripalani2007deficits]. The success of large language models (LLMs) has provided a promising avenue for this clinically important task of BHC synthesis. Having LLMs generate the first draft of a BHC may lead to fewer errors, higher-quality summaries, and reduced clinician time spent on documentation.
@@
Recent advancements in natural language processing (NLP) have been characterized by transformer-based language models [NIPS2017_3f5ee243], such as Bidirectional Encoder Representations from Transformers (BERT) [kenton2019bert] and Unidirectional Generative Pre-trained Transformer 2 (GPT-2) [radford2019language]. The evolution of transformer models involved pretraining on large corpora of text and subsequent fine-tuning on domain-specific data, as demonstrated by GPT-3 [NEURIPS2020_1457c0d6], PaLM [chowdhery2023palm], and T5 [raffel2020exploring]. After pre-training, these LLMs are often exposed to supervised fine-tuning given input-output pairs to improve their ability to follow instructions (instruction turning) [zhang2023instruction]. Following instruction tuning, reinforcement learning with human feedback (RLHF) is employed to align the LLM with human preferences [zhang2023instruction, wang2023pre]. A recent work presented a multi-document brief hospital course dataset to promote the advancement of hospital visit summarization [adams2021s]. Furthermore, deep learning-based automation methods for hospital course summarization have been explored [searle2023discharge, hartman2023method], but none provide an open-source benchmark for state-of-the-art (SOTA) LLMs.
We utilize the MIMIC-IV-Note [johnson2023mimic] dataset to curate a foundational and clinically relevant dataset, the MIMIC-IV-BHC [aali2024mimic]. Subsequently, we present an open-source benchmark of LLM performance on this dataset. We compare SOTA LLMs: proprietary general-purpose LLMs, such as GPT-3.5 [openai2022chatgpt] and GPT-4 [bubeck2023sparks, openai2023gpt4], open-source LLMs, such as Llama2 [touvron2023llama], FLAN- UL2 [tay2022unifying], and a T-5 model pre-trained on medical text [lehman2023clinical]. We explore prompting-based (in-context learning [ICL]) and fine-tuning-based adaptation strategies for our benchmarking study’s proprietary and open-source LLMs to identify the best model and adaptation strategy. We employ well-established metrics that measure syntactic and semantic similarity between LLM-generated and clinician-written BHCs. Our analysis allows us to explore the robustness of domain-adapted LLMs in handling BHCs derived from clinical notes with varying context lengths. Finally, we conduct a rigorous clinical reader study involving five expert clinicians reviewing clinical notes and BHC pairs to ensure adequate qualitative evaluation of the LLM-generated summaries across dimensions of comprehensiveness, conciseness, factual correctness, and fluency.
@@
We utilize a raw clinical data source [albers2020assimilation], the MIMIC-IV-Note dataset [johnson2023mimic], a compilation of 331,794 de-identified discharge summaries from 145,915 patients admitted to the Beth Israel Deaconess Medical Center. MIMIC-IV-Note [johnson2023mimic] is a publicly available clinical dataset reflective of real-world clinical notes. We pass MIMIC-IV [johnson2023mimic] notes through our pre-processing pipeline to create a labeled task-specific dataset, the MIMIC-IV-BHC [aali2024mimic] (Table LABEL:table:dataset_sample), emphasizing the relationship between clinical notes and BHCs. Our open-source novel dataset is crucial to allow researchers to pursue further benchmarking studies.
@@
We perform data translation [albers2020assimilation] by pre-processing raw clinical notes and taking key steps to ensure the dataset’s standardization and suitability for downstream machine-learning tasks. Using regular expression (RegEx) and natural language toolkit (NLTK) in Python, we apply: whitespace removal, uniform spacing, section identification, header standardization, extraneous symbol removal, and tokenization. We extract the BHC from each clinical note to create the labeled MIMIC-IV-BHC [aali2024mimic] (Table LABEL:table:dataset_sample) dataset containing 270,033 clinical notes–BHC pairs with an average input token length of 2,267 $\pm$ 914 and an average output token length of 564 $\pm$ 410. We apply a uniform set of processing steps across all samples and manually review 100 randomly drawn samples from our dataset to validate the effectiveness of the pre-processing pipeline. Our clinicians manually review 30 clinical note-summary pairs from the dataset, without reporting any quality issues (Figure 1).
@@
  • ChatGPT (API version GPT-3.5-Turbo) [openai2022chatgpt] contains 175 billion parameters and has been extensively fine-tuned for general tasks using human feedback, enhancing its ability to capture intricate details within any summarization task. GPT-3.5 can support input context lengths of up to 16,384 tokens.
  • GPT-4 (API version GPT-4-0613) [openai2023gpt4] achieves SOTA NLP performance and supports a context window of 128,000 tokens, the highest compared to models listed in Figure 2, making it a suitable choice in multi-document summarization scenarios.
@@
To identify the best-performing open-source and proprietary models, we conduct a performance analysis, where we select a subset of 70 independent samples from the 100 test samples in the 0-1,024 context range. The remaining 30 independent samples from this range were kept separate for qualitative evaluation. This distinct evaluation set helps mitigate data leakage, providing a more objective assessment of model generalization while ensuring that the samples used for quantitative evaluation are not included in the reader study. To accurately assess the performance of models, we employ quantitative metrics commonly used in summarization tasks to evaluate the syntactic similarity and semantic relevance of the summaries. Bilingual Evaluation Understudy (BLEU) [papineni2002bleu] evaluates the overlap between the generated and reference summaries using a weighted average of 1 to 4-gram precision. ROUGE-L [lin2004rouge] assesses the precision and recall of the longest common subsequence. BLEU and ROUGE-L depict lexical overlap between generated and reference summaries. In contrast, we also use semantic similarity metrics like BERT-Score [zhang2019bertscore] which leverages contextual BERT embeddings to evaluate the similarity between the generated and reference summaries. Finally, we perform an independent context-length analysis using fine-tuned Llama2-13B and GPT-4 with ICL, the best-performing LLMs allowing context lengths greater than 4,000 tokens as input using 100 test samples from each of the three context-length test sets from Table LABEL:table:dataset_sampleb. To assess potential health equity implications, we perform a stratified subgroup analysis across patient-reported sex when assessing the quantitative performance of BHC summarization techniques.
@@
As part of the study, we request optional qualitative feedback from clinicians on each note-BHC pair to dive deeper into the reasons behind clinician preferences in the four initial categories: 1) Factual Mistakes: Does the summary inaccurately represent critical information from the original note?, 2) Missing Critical Information: Does the summary omit critical information part of the original note?, 3) Hallucinations: Does the summary contain information not present in the original note?
@@ RESULTS
We find that our five clinical readers strongly prefer in-context GPT-4 summaries for attributes of comprehensiveness, conciseness, factual correctness, and fluency, compared to the original clinician-written summaries ($p<0.001$) and QLoRA Llama2-13B summaries ($p<0.001$) (Figure 5a). However, we find no significant differences between reader preferences comparing Llama2 summaries and the original clinician-written summaries ($p=0.05$). Overall, GPT-4 scores have a narrower range, showcasing lower variation. We report the proportion of scores from readers between the range of 3 to 5 for GPT-4, Llama2-13B, and clinician summaries: 100%, 47%, and 40%, respectively. To further explore reasons behind clinician preferences, we analyze 37 optional "free-text" comments provided by clinicians and report common themes with respective frequencies in Figure 5b. In-context GPT-4 summaries consistently report the least amount of cases across all themes: factual mistakes, omission of critical information, and hallucinations. QLoRA Llama2-13B and clinician-written summaries exhibit similar frequencies across all themes. This study underscores the strength of LLMs in generating summaries that are both statistically superior and preferred by our panel of clinicians in the majority of cases.
@@ DISCUSSION
Recent adaptations of LLMs in the clinical domain emphasize their potential in understanding medical language. Noteworthy recent works [van2023radadapt, van2023clinical, chen2022toward, lyu2023translating, patel2023chatgpt, singh2023chatgpt] use open-source datasets and display the performance of LLMs in multiple summarization tasks in medicine: radiology reports, patient health questions, progress notes, doctor-patient conversations, and discharge summaries. Other works [wang2024drg, singhal2023large] further explore domain adaptation of LLMs in the clinical domain through fine-tuning for improved performance in medical tasks. Our benchmark shows that open-source LLMs can produce high-quality BHCs and that Llama2-13B comes out as a superior choice for further analysis of open-source LLMs based on its performance and an expanded context length window (Figure 3a). The subgroup analysis (Figure 3b) further emphasizes the importance of a multi-faceted analysis when evaluating models. Furthermore, our context-length analysis (Figure 4) showcases the benefit of fine-tuning LLMs on datasets with a wide range of context lengths. A deeper qualitative analysis through our reader study (Figure 5) suggests that adapted proprietary LLMs like GPT-4 can outperform other adapted LLMs as well as clinicians, while open-source LLMs like Llama2-13B can produce summaries similar in quality to clinician summaries. However, open-source models provide the benefit of ease in implementation, where sending data via an API may not be possible due to privacy constraints. Our benchmark also reports strong summarization performance from Clinical-T5-Large, while it is the smallest parameter model, possibly due to extensive pre-training on medical text. However, Clinical-T5-Large also presents a large variation in performance across subgroups, making it less suitable.
@@
For our quantitative benchmarking study, we employ a targeted evaluation strategy by including a slew of metrics whose individual strengths cover the weaknesses of other metrics. We reported ROUGE-L, BERT-Score, and BLEU for shorter context length pairs (0-1,024 tokens) as these metrics remain widely used for evaluating summaries in this length regime. We report BERT-Score for longer context length pairs ($>1,024$ tokens) recognizing the potential shortcomings of ROUGE-L for longer summaries and because BERT-Score offers a more semantic similarity assessment.
@@
Our study bridges computational and clinical disciplines by combining quantitative NLP metrics with qualitative clinical assessments, ensuring the evaluation reflects both technical performance and clinical relevance. It is crucial to recognize that quantitative measures, such as NLP metrics, and qualitative assessments by clinical readers gauge different aspects of summarization quality. NLP metrics quantify the level of (textual, semantic, contextual) similarity between two summaries, whereas clinical readers inform perceptual (subjective) quality differences. We observe that fine-tuned Llama2-13B consistently outperforms in-context GPT-4 in quantitative similarity metrics. However, the reader study results present a nuanced narrative where clinicians express a preference for GPT-4-generated summaries. This finding reflects a critical point: high quantitative similarity scores do not necessarily translate to a "better summary" from a clinician’s perspective, as well as across subgroups. Enhancing the alignment between quantitative metrics and qualitative assessments will enable a more holistic evaluation for clinical text summarization [van2023clinical].
@@
While categorizing models as strictly open-source or proprietary may not comprehensively capture the many differences between the models, we opt for this categorization because fine-tuning proprietary models is possible but only with unreported mechanisms, whereas, fine-tuning open-source LLMs can lead to repeatable and verifiable models. Further, within each category, we justify the model choice for the broader quantitative study by aiming for diverse models, considering factors like model accessibility and architecture. We fine-tune only a subset of models containing up to 20 billion parameters (such as Llama2-13B, Clinical-T5-Large, and FLAN-UL2), as they can commonly be fine-tuned on consumer-grade GPUs. In contrast, for the qualitative study, we select the best model from each of the open-source and proprietary categories based on their performance in the quantitative study.
@@
While noting the strengths of our study, we also discuss its limitations. The first limitation we observe is that because Clinical-T5-Large can only accept inputs of less than 512 tokens, the quality of the Clinical-T5-Large generated summaries in our experiments might be impacted. However, recent works [van2023radadapt] have demonstrated the strong performance of domain-adapted seq2seq models from the T5 family in radiology report summarization, motivating our inclusion of this model. Moreover, based on prior work [fleming2024medalign], given strong decoder models, there is potential to use multi-stage refinement and summarization techniques to condense the size of the input prompt itself. Another limitation of our work is the limited availability of publicly available datasets for clinical note summarization, making it difficult to gauge the performance of our adapted LLMs in diverse scenarios across hospitals and clinical practices. A final limitation of our work is that adaptation strategies and larger LLMs released after our experimentation phase could not be explored in the study. However, larger models often demonstrate superior performance in summarization tasks, and our existing framework, dataset, and software can easily be used to explore newer LLM families. Hence, we motivate future studies to iterate on newer models and adaptation strategies given the MIMIC-IV-BHC [aali2024mimic] dataset.

### v3

@@ Abstract
Materials and Methods: Using clinical notes as input, we apply prompting-based (using in-context learning) and fine-tuning-based adaptation strategies to three open-source LLMs (Clinical-T5-Large, Llama2-13B, FLAN-UL2) and two proprietary LLMs (GPT-3.5, GPT-4). We evaluate these LLMs across multiple context-length inputs using natural language similarity metrics. We further conduct a clinical study with five clinicians, comparing clinician-written and LLM-generated BHCs across 30 samples, focusing on their potential to enhance clinical decision-making through improved summary quality. We compare reader preferences for the original and LLM-generated summary using Wilcoxon Signed-Rank tests. We further request optional qualitative feedback from clinicians to gain deeper insights into their preferences, and we present the frequency of common themes arising from these comments.
@@
Clinicians spend significant time on clinical documentation [moy2021measurement, chaiyachati2019assessment, mamykina2016residents], a crucial component of data assimilation [albers2020assimilation] beginning with data sources such as EHRs and progressing through data translation and knowledge creation where these records can be processed for clinically-relevant machine learning tasks. The discharge summary (clinical note) is a key document clinicians refer to in important scenarios, including when a discharged patient presents to the clinic for post-hospitalization follow-up. The brief hospital course (BHC) fulfills an important role within the discharge summary, particularly for long and complex hospitalizations. Clinicians who are pressed for time may choose to focus on the BHC summary instead of the entire discharge summary. Synthesizing a BHC given a clinical note requires substantial clinician time and expertise, with errors possibly causing patient harm [clough2024transforming]. Automating BHC generation is important since discharge summaries can lack essential information, contain incorrect information, and may be incomplete before follow-up appointments [kripalani2007deficits]. The success of large language models (LLMs) has provided a promising avenue for this clinically important task of BHC synthesis. Having LLMs generate the first draft of a BHC may lead to fewer errors by reducing cognitive load and fatigue-related mistakes, produce higher-quality summaries by consistently incorporating all relevant clinical details, and substantially reduce clinician time spent on documentation by automating the initial drafting process [patel2023chatgpt, singhal2023large, van2024clinical]. However, while LLMs can produce high-quality drafts, clinical use should consider automation bias, where these drafts could be accepted as final without a thorough review, requiring interfaces designed to promote critical assessment and evaluation [warraich2024fda].
@@
Recent advancements in natural language processing (NLP) have been characterized by transformer-based language models [NIPS2017_3f5ee243], such as Bidirectional Encoder Representations from Transformers (BERT) [kenton2019bert] and Unidirectional Generative Pre-trained Transformer 2 (GPT-2) [radford2019language]. The evolution of transformer models involved pretraining on large corpora of text and subsequent fine-tuning on domain-specific data, as demonstrated by GPT-3 [NEURIPS2020_1457c0d6], PaLM [chowdhery2023palm], and T5 [raffel2020exploring]. After pre-training, these LLMs are often exposed to supervised fine-tuning given input-output pairs to improve their ability to follow instructions (instruction turning) [zhang2023instruction]. Following instruction tuning, reinforcement learning with human feedback (RLHF) is employed to align the LLM with human preferences [zhang2023instruction, wang2023pre]. A recent work presented a multi-document brief hospital course dataset to promote the advancement of hospital visit summarization [adams2021s].
Automating documentation tasks such as BHC synthesis can potentially alleviate clinician workload and improve documentation quality. While previous works have explored deep learning-based automation for hospital course summarization [searle2023discharge, hartman2023method, jung2024enhancing], no open-source benchmark exists to evaluate state-of-the-art (SOTA) LLM performance on BHC synthesis. Developing a benchmark and standardized framework for evaluating model effectiveness can help advance NLP applications in reducing clinical documentation burdens. Our study addresses this gap by introducing a foundational task-specific dataset, MIMIC-IV-BHC [aali2024mimic], which is tailored specifically for BHC summarization. By benchmarking SOTA LLMs on this dataset, we aim to facilitate the development of more efficient and accurate models for brief hospital course summarization. The significance of this work lies in its potential to reduce documentation errors, improve the quality of clinical summaries, and free up clinician time, which could directly contribute to enhancing patient care and safety.
@@
We utilize the MIMIC-IV-Note [johnson2023mimic] dataset to curate a foundational task-specific dataset, the MIMIC-IV-BHC [aali2024mimic], tailored for the task of BHC summarization. Subsequently, we present an open-source benchmark of LLM performance on this dataset, across a variety of adaptation strategies.
@@
We utilize a raw clinical data source [albers2020assimilation] called MIMIC-IV-Note [johnson2023mimic], a compilation of 331,794 de-identified discharge summaries from 145,915 patients admitted to the Beth Israel Deaconess Medical Center. MIMIC-IV-Note [johnson2023mimic] is a publicly available dataset reflective of real-world clinical notes. We pre-process MIMIC-IV [johnson2023mimic] notes to create a labeled task-specific dataset, the MIMIC-IV-BHC [aali2024mimic] (Table LABEL:table:dataset_sample), emphasizing the relationship between clinical notes and BHCs. Our open-source novel dataset is crucial to allow researchers to pursue further benchmarking studies. This study was institutional review board exempt because the MIMIC dataset does not meet the criteria for human subjects research described by our institution.
@@
We perform data translation [albers2020assimilation], a process of transforming raw, unstructured data into a standardized and structured format for downstream machine-learning tasks. This includes steps such as whitespace removal, section identification, tokenization, and other transformations that create a uniform structure for the MIMIC-IV-BHC dataset [aali2024mimic]. Using regular expressions and natural language toolkit (NLTK) in Python, we apply: whitespace removal, uniform spacing, section identification, header standardization, extraneous symbol removal, and tokenization. We extract the BHC from each clinical note to create the labeled MIMIC-IV-BHC [aali2024mimic] (Table LABEL:table:dataset_sample) dataset containing 270,033 clinical notes–BHC pairs with an average input token length of 2,267 $\pm$ 914 and an average output token length of 564 $\pm$ 410. We apply a uniform set of processing steps across all samples and manually review 100 randomly drawn samples from our dataset to validate the effectiveness of the pre-processing pipeline. The choice of 100 samples was based on a balance between confirming data quality and the resources available for manual review. Furthermore, the preprocessed dataset underwent an extensive review process before its publication on PhysioNet. Finally, our clinicians manually reviewed 30 clinical note-summary pairs from the dataset, without reporting any quality issues (Figure 1).
@@
  • GPT-3.5 (Azure API version GPT-3.5-Turbo-0613) [openai2022chatgpt] contains 175 billion parameters and has been extensively fine-tuned for general tasks using human feedback, enhancing its ability to capture intricate details within any summarization task. This GPT-3.5 version can support input context lengths of up to 16,384 tokens.
  • GPT-4 (Azure API version GPT-4-0613) [openai2023gpt4] achieves SOTA NLP performance and supports a context window of 32,768 (later expanded to 128,000) tokens, the highest compared to models listed in Figure 2, making it a suitable choice in multi-document summarization scenarios.
@@
To identify the best-performing open-source and proprietary models, we conduct a performance analysis by randomly selecting 70 independent samples from the 100 test samples in the 0-1,024 context range. We reserve the remaining 30 independent samples for qualitative evaluation by clinicians. The 70:30 split between quantitative and qualitative analysis was determined to ensure a significant portion of the dataset is used for quantitative evaluation while keeping the workload for clinicians manageable. This separation also mitigates data leakage, providing a more objective assessment of model generalization, as the samples used for quantitative evaluation were not included in the reader study. To accurately assess the performance of models, we employ quantitative metrics commonly used in summarization tasks to evaluate the syntactic similarity and semantic relevance of the summaries. Bilingual Evaluation Understudy (BLEU) [papineni2002bleu] evaluates the overlap between the generated and reference summaries using a weighted average of 1 to 4-gram precision. ROUGE-L [lin2004rouge] assesses the precision and recall of the longest common subsequence. We focus only on ROUGE-L instead of other ROUGE variants to avoid redundancy, as it effectively captures essential n-gram overlap in our context and correlates well with other ROUGE variants [liu2009exploring]. BLEU and ROUGE-L depict lexical overlap between generated and reference summaries. In contrast, we also use semantic similarity metrics like BERT-Score [zhang2019bertscore] which leverages contextual BERT embeddings to evaluate the similarity between the generated and reference summaries. Finally, we perform an independent context-length analysis using fine-tuned Llama2-13B and GPT-4 with ICL, the best-performing LLMs allowing context lengths greater than 4,000 tokens as input using 100 test samples from each of the three context-length test sets from Table LABEL:table:dataset_sampleb. Output lengths were set to the models’ default configurations and were not specifically controlled in this analysis. To assess potential health equity implications, we perform a stratified subgroup analysis across patient-reported sex when assessing the quantitative performance of BHC summarization techniques.
@@
For the reader study, we gather diverse and well-informed feedback by ensuring that all participating clinicians possess at least five years of experience and a solid understanding of discharge summaries—a standard yet crucial task for clinicians. To reduce potential bias, we randomize the order of the three summaries, labeling them as ’Summary 1,’ ’Summary 2,’ or ’Summary 3.’ All clinicians see questions in the same order, helping maintain consistency and reduce potential biases. To standardize score values, we provide a detailed rating scale across each aspect. For example, in rating comprehensiveness, a score of 1 indicates minimal capture of critical information, whereas a score of 5 denotes full coverage of key details. Furthermore, we guide clinicians to comparatively assess summaries within each note, ensuring that the scores consistently reflect differences in quality. To quantitatively evaluate the consistency of ratings, we calculate Krippendorff’s alpha based on all scores provided by clinicians.
In total, we received 150 reviews across all five clinicians (30 clinical note-summary pairs for each of the five clinicians). As part of the study, we request optional qualitative feedback from clinicians on each note-BHC pair to dive deeper into the reasons behind clinician preferences. Of the 150 reviews, 37 received optional free-text comments from clinicians, while the others were left blank. We analyze all 37 comments to determine insights from observed common themes: 1) Factual Mistakes: Does the summary inaccurately represent critical information from the original note?, 2) Missing Critical Information: Does the summary omit critical information part of the original note?, 3) Hallucinations: Does the summary contain information that cannot be inferred from the original note?
@@ RESULTS
We find that our five clinical readers strongly prefer in-context GPT-4 summaries for attributes of comprehensiveness, conciseness, factual correctness, and fluency, compared to the original clinician-written summaries ($p<0.001$) and QLoRA Llama2-13B summaries ($p<0.001$) (Figure 5a). However, we find no significant differences between reader preferences comparing Llama2 summaries and the original clinician-written summaries ($p=0.05$). Overall, GPT-4 scores have a narrower range, showcasing lower variation. We report the proportion of scores from readers between the range of 3 to 5 for GPT-4, Llama2-13B, and clinician summaries: 100%, 47%, and 40%, respectively. To further explore reasons behind clinician preferences, we analyze 37 optional "free-text" comments provided by clinicians and report common themes with respective frequencies in Figure 5b. In-context GPT-4 summaries consistently report the least amount of cases across all themes: factual mistakes, omission of critical information, and hallucinations. QLoRA Llama2-13B and clinician-written summaries exhibit similar frequencies across all themes. As part of the reader study, we observe moderate consistency among raters given the inherently subjective nature of the task (Krippendorff’s $\alpha=0.47$). Overall, this study underscores the strength of LLMs in generating summaries that are both statistically superior and preferred by our panel of clinicians in the majority of cases.
@@ DISCUSSION
Recent adaptations of LLMs in the clinical domain emphasize their potential in understanding medical language. Noteworthy recent works [van2023radadapt, van2024clinical, chen2022toward, lyu2023translating, patel2023chatgpt, singh2023chatgpt] use open-source datasets and display the performance of LLMs in multiple summarization tasks in medicine: radiology reports, patient health questions, progress notes, doctor-patient conversations, and discharge summaries. Other works [wang2024drg, singhal2023large] further explore domain adaptation of LLMs in the clinical domain through fine-tuning for improved performance in medical tasks. Our benchmark shows that open-source LLMs can produce high-quality BHCs and that Llama2-13B comes out as a superior choice for further analysis of open-source LLMs based on its performance and an expanded context length window (Figure 3a). The subgroup analysis (Figure 3b) further emphasizes the importance of a multi-faceted analysis when evaluating models. Furthermore, our context-length analysis (Figure 4) showcases the benefit of fine-tuning LLMs on datasets with a wide range of context lengths. A deeper qualitative analysis through our reader study (Figure 5) suggests that adapted proprietary LLMs like GPT-4 can outperform other adapted LLMs as well as clinicians, while open-source LLMs like Llama2-13B can produce summaries similar in quality to clinician summaries. However, open-source models provide the benefit of ease in implementation, where sending data via an API may not be possible due to privacy constraints. Our benchmark also reports strong summarization performance from Clinical-T5-Large, while it is the smallest parameter model, possibly due to extensive pre-training on medical text. However, Clinical-T5-Large also presents a large variation in performance across subgroups, making it less suitable.
@@
For our quantitative benchmarking study, we employ a targeted evaluation strategy by including a slew of metrics whose individual strengths cover the weaknesses of other metrics. We reported ROUGE-L, BERT-Score, and BLEU for shorter context length pairs (0-1,024 tokens) as these metrics remain widely used for evaluating summaries in this length regime. For longer context length pairs (>1,024 tokens), we report only BERT-Score, recognizing the potential shortcomings of ROUGE-L for longer summaries. ROUGE-L primarily captures syntactic matches rather than semantic similarity, which can result in limited content coverage and a preference for summaries that closely follow the original wording [lin2004rouge, zhang2019bertscore, koh2022far, van2024clinical]. BERT-Score offers a more semantic similarity assessment, though it too may underrepresent fine-grained information overlap, particularly in long-form texts [zhang2019bertscore]. We also acknowledge that there is currently no perfect metric for evaluating complex, long-form summarizations in clinical settings [van2024clinical].
@@
Our study bridges computational and clinical disciplines by combining quantitative NLP metrics with qualitative clinical assessments, ensuring the evaluation reflects both technical performance and clinical relevance. It is crucial to recognize that quantitative measures, such as NLP metrics, and qualitative assessments by clinical readers gauge different aspects of summarization quality. NLP metrics quantify the level of (textual, semantic, contextual) similarity between two summaries, whereas clinical readers inform perceptual (subjective) quality differences. We observe that fine-tuned Llama2-13B consistently outperforms in-context GPT-4 in quantitative similarity metrics. However, the reader study results present a nuanced narrative where clinicians express a preference for GPT-4-generated summaries. This finding reflects a critical point: high quantitative similarity scores do not necessarily translate to a "better summary" from a clinician’s perspective, as well as across subgroups. Enhancing the alignment between quantitative metrics and qualitative assessments will enable a more holistic evaluation for clinical text summarization [van2024clinical].
@@
While categorizing models as strictly open-source or proprietary may not comprehensively capture the many differences between the models, we opt for this categorization because fine-tuning proprietary models is possible but only with unreported mechanisms, whereas, fine-tuning open-source LLMs can lead to repeatable and verifiable models. For the broader quantitative study, we justify the model choice by aiming for diverse models, considering factors like model accessibility and architecture. In contrast, for the qualitative study, we select the best model from each of the open-source and proprietary categories based on their performance in the quantitative study. From the proprietary LLM category, we select GPT-4 (SOTA at the time of experimentation), serving as a robust reference for proprietary models, and GPT-3.5, a smaller, more cost-effective comparison within the same category. Due to resource constraints and the proprietary nature of fine-tuning mechanisms for GPT-3.5 and GPT-4, we limit fine-tuning to open-source models with up to 20 billion parameters (such as Llama2-13B, Clinical-T5-Large, and FLAN-UL2), as these can be fine-tuned on consumer-grade GPUs.
@@
While noting the strengths of our study, we also discuss its limitations. The first limitation we observe is that because Clinical-T5-Large can only accept inputs of less than 512 tokens, the quality of the Clinical-T5-Large generated summaries in our experiments might be impacted. However, recent works [van2023radadapt] have demonstrated the strong performance of domain-adapted seq2seq models from the T5 family in radiology report summarization, motivating our inclusion of this model. Moreover, based on prior work [fleming2024medalign], given strong decoder models, there is potential to use multi-stage refinement and summarization techniques to condense the size of the input prompt itself. Another limitation of our work is the limited availability of publicly available datasets for clinical note summarization, making it difficult to gauge the performance of our adapted LLMs in diverse scenarios across hospitals and clinical practices. The MIMIC-IV dataset is derived from a specific U.S.-based healthcare system, and its generalizability to other healthcare systems worldwide may be limited. Similarly, the clinician population in our study may not be fully representative of broader clinical populations. Additional public clinical note datasets may help alleviate these challenges. We also acknowledge that our clinician reader mix, primarily composed of radiology subspecialists with one clinical physician, may affect the generalizability of findings, as discharge summaries are typically reviewed by clinical doctors. A final limitation of our work is that newer adaptation strategies and LLMs, such as Llama-3 (released after our experimentation phase), could not be incorporated into this study. However, our framework, and dataset are designed to be model agnostic, enabling future studies to evaluate newer LLM families and adaptation strategies as they emerge. We encourage further research to iterate on these advancements using the MIMIC-IV-BHC dataset [aali2024mimic].